OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) <= 25?

OpenStudy (lalaly):

the following?

OpenStudy (anonymous):

x >= -3 x <= -3 -3 <= x A number line with a closed circle on negative 3 and shading to the right

OpenStudy (lalaly):

second one

OpenStudy (radar):

7-9x-x-12<=25 -10x<=25+5 -10x<=30 x>=-3 Don't forget when a division or multiplication by a negative you reverse the sense of the equality sign (direction)

OpenStudy (radar):

That is why it is the first one. It should be noted that the first one and the third one express the same thing. x is equal to or larger than -3 !!